A processor of carrots cuts the green top off each carrot, washes the carrots, and inserts six to a package. Twenty packages are inserted in a box for shipment. To test the mass of the boxes, a few were checked. The mean mass was 9.2 kg, the standard deviation 0.27 kg. How many boxes must the processor sample to be 99% confident that the sample mean does not differ from the population mean by more than 0.14 kg? (Round the intermediate calculation to 2 decimal places. Round the final answer to the nearest whole number.)

Sample size      

Exercise 8-37 Suppose the prime minister wants an estimate of the proportion of the population who support his current policy    
Answer #1

Q.1) Given that, population standard deviation (\sigma) = 0.27 kg

Margin of error ( E ) = 0.14 kg

A 99% confidence level has significance level of 0.01 and critical value is, Zo/2 = 2.575

We want to find, the sample size ( n ),

0.27 0.14 ->nN 25

Therefore, required sample size is 25

Q.2) Given that, margin of error ( E ) = 0.08

A 80% confidence level has significance level of 0.20 and critical value is, a21.28

We want to find, sample size for following cases,

a) If estimate of proportion = p = 0.35

1.282 n = p * (1-p) * (4y2)2 = 0.35 * (1-0.35) * (しの)-58.24 0.08 > n 58

Therefore, required sample size is 58

b) If no estimate is given then we assume p = 0.5

n=ps(1-p) * (-_リー)2-0.35 * (1-0.5) * ( -)2-64 0.08 64

Therefore, required sample size is 64